OpenStudy (anonymous):

what is tan(3pi/4)? Explain your thinking.

OpenStudy (anonymous):

so you want to know what it is? right?

OpenStudy (anonymous):

\[\tan \frac{ 3 \pi }{ 4 }=\tan \left( \pi-\frac{ \pi }{ 4 } \right)=-\tan \frac{ \pi }{ 4 }=?\]

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

so that's the answer.... just -1?